A block of mass 4kg is on the point of slipping down a plane inclined at an angle of 30° to the horizontal.What force parallel to the plane will just move it up the plane?


1
Expert's answer
2022-02-10T14:02:27-0500

Solution

Inclined angle

"\theta=30°"

Mass m=4Kg



force parallel to the plane will just move it up the plane is given by

F=mg sin "\theta"

"=4\times9.81\times sin30°\\=19.62N"